OpenStudy (aye_my_peeps):

a conveyer belt moves 450ft in 3/8 hr. complete the unit rate

OpenStudy (mahoganie.carson):

3/8 = 0.375 (h) f=450 ft/h 450/0.375=1200

OpenStudy (aye_my_peeps):

so its 1200\

OpenStudy (mahoganie.carson):

Yes it is 1200 :)